Hi I am trying to solve this system
2x+4y=8
3x+10y=18
I tried the addition method and came up with this:
3(2x+4y)=3(8)-
2(3x+10y)= 2(18)=
-8y=-12=
y=2/3
but I know this is not the right answer because the original problem was in dollars and you can't have 2/3rds of a dollar

Answer by mananth(16949)

2x+4y=8 .............1
3x+10y=18 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 2
-10x-20y=-40
6x+20y=36
Add the two equations
-4x=-4
/-4
x=1.00
plug value of x in (1)
2x+4y=8
2+4y=8
4y=8-2
4y=6
y=1.50